OpenStudy (anonymous):

Skylar's grades on four math tests are 85, 78, 77, and 69. What does Skylar need to score on the next test in order to have a mean score of 80? 83 90 91

Parth (parthkohli):

Do you know how to solve equations?

Parth (parthkohli):

If you need a hint, just mention me in the replies.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow {x + 85 + 78 + 77 + 69 \over 5} = 80 }\) Challenge to solve for x, where x will be the answer!

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

91

Parth (parthkohli):

You have to find the value of x in the equation. It's because x is a number we don't know so we can put it as a variable. Now, if you know the formula, then you know why put that equation.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x + 85 + 78 + 77 + 69 = 400 }\) \(\Large \color{purple}{\rightarrow x + 309 = 400 }\) \(\Large \color{purple}{\rightarrow x = 91 }\) YES!